# Diffusion models for multivariate subsurface generation and efficient probabilistic inversion

### v1

@@ 2.5 Code implementation
The methods were implemented in Python (Python 3.12) using PyTorch (PyTorch 2.5) libraries with CUDA implementation. The training of all the generative models and the corresponding unconditional and conditional modeling were performed using a single GPU NVIDIA GeForce® GTX TITAN X (CUDA 12.2). The pCN sampling was performed on the same machine, generating DM samples using this GPU. No parallelization on different GPUs was considered. Training, modeling and sampling times mentioned in Section 3.3 are related to this machine.
@@ 3 Results
We evaluate the generative performance of the trained DM by parameterizing a bivariate prior pdf, describing a sequence of lenticular sandy channels in a shale background (facies), and the associated acoustic impedance ($I_{P}$). We consider an area of 100 $\times$ 80 cells, assumed to have sides of 1 m (Fig. 1). The TI consists of 3000 samples (Fig. 1a) of MPS facies realizations, generated using Petrel software (SLB) based on a 3D conceptual geological model. Shales and sands are categorized as 0 and 1, respectively. The facies distribution ensemble is spatially uniform, with each cell having average p(sand) = 0.3 (Fig. 1c). The facies-dependent $I_{P}$ is represented using direct sequential co-simulations with multi-local distribution functions (Nunes et al., 2017; Soares, 2001). We account for facies-dependent $I_{P}$ marginal distributions (Fig. 1b) and spatial models (2D variograms; detailed in Appendix A). The average spatial distribution of $I_{P}$ (Fig. 1d) is primarily dependent on facies with an area of relatively lower values in the lower-right section of the considered area. The DM was trained for a total of 600 epochs (225000 steps). The GAN and VAE used for benchmarking were trained on the same TI, for a total of 2000 epochs, using the hyperparameters proposed by Miele et al. (2024).
Figure 1: Prior distribution as represented in the training dataset. a) Facies and $I_{P}$ realizations; b) $I_{P}$ distribution per facies; c) point-wise average facies distribution; d) point-wise average and standard deviation of $I_{P}$.
@@ 3.1 Unconditonal modeling
Prior modeling performances are assessed on 1000 unconditional samples generated from the trained DM, GAN and VAE each. The channel morphologies and $I_{P}$ spatial patterns reproduced by DM and GAN (Samples in Fig. 2a and c) are visually indistinguishable from those of the TI (Fig. 1a). For DM, the p(sand) and average $I_{P}$ match the prior accurately; Table 1 shows the average relative error for the two property types. Of the two benchmarking models, the GAN’s p(sand) and average $I_{P}$ distributions (Fig. 2c) show significant misfit with the prior. In fact, the overestimation of p(sand) in the central area may indicate mode collapse. Despite the smooth spatial features in the generated samples (Fig. 2e), the VAE performs better than GAN at capturing both properties average distributions, and displays even a larger accuracy than DM for facies (Table 1).
The joint distributions of facies and co-located $I_{P}$ obtained from the sampled realizations (Fig. 2b, d, and f) show significantly superior modeling ability of the DM (Fig. 2b). The facies classes are modeled with negligible errors (i.e., facies values different from 0 and 1); the corresponding $I_{P}$ marginal distributions (plotted along the y-axes) match accurately the prior’s mean and standard deviation values, fitting the data with a $KL$ divergence approximately five times lower than for GAN and VAE (Table 1). Further analysis of the networks unconditional modeling performances focusing on geostatistical metrics, are provided in Appendix A.
Table 1: Metrics for unconditional prior modeling using the trained DM and benchmarking GAN and VAE. The p(sand) and Avg. $I_{P}$ refer to the average models obtained from the realizations ensembles, shown in Figure 2a, c, e. $KL$: Kullback–Leibler divergence; Avg.: average; St. dev.: standard deviation.
@@
7.0%
1.2%
6870 ±760
8620 ±690
1.3e-3
1.5e-3
Figure 2: Summary of unconditional modeling performances of DM (a, b), GAN (c, d), and VAE (e, f). Subplots (a), (c), (e): realizations and average distribution of facies and $I_{P}$ for DM (a), GAN (c) and VAE (e); subplots (b), (d), (f): joint and $I_{P}$ distributions for DM (b), GAN (d) and VAE (f).
@@ 3.2 Inverse modeling using CDPS
We evaluated the proposed CDPS by considering both linear and nonlinear inverse problems using a test realization from the prior distribution as our "True" facies and $I_{P}$ (Fig. 3a). The linear test case consists in generating samples that are locally conditioned on direct observations. The data, $\bf{d_{obs}}$, are obtained from two wells ("Well 1" and "Well 2" in Figure 3a). This task is analogous to image inpainting in computer vision. The nonlinear example is a geophysical inversion problem with fullstack seismic data (Fig. 3c), assumed to be derived from the true $I_{P}$, and modeled as follows. First, vertical sequences of impedance contrasts ($R$) are computed following $R_{i}=(I_{P_{i+1}}-I_{P_{i}})/(I_{P_{i+1}}+I_{P_{i}})$, where the subscript $i$ indicates an $I_{P}$ sample. Then, each trace is convolved with a known seismic source wavelet; in our case, a Ricker wavelet with central frequency of 25Hz and resolution of 3 ms. For both case studies, we evaluated the CDPS under different Gaussian data noise conditions (Table 2).
@@
Using 300 validation examples from the TI we evaluated the modeling error $\sigma_{\hat{\textbf{x}}_{0}}^{(t)}$ for 1000 denoising steps. Examples of pixel-dependent uncertainty are shown in Fig. 4a, the full homoscedastic noise sequence used is shown in Fig. 4b. Here, even if the facies classes are categorical, they are modeled as a continuous variable by DM and treated as such in the likelihood score computation. In the Appendix B we applications of CDPS using both well and seismic data, simultaneously.
Figure 3: Test model used in inverse modeling; a) real subsurface distribution ("True") and location of the two wells used for the linear inversion; b) conditioning well log data for the linear inversion; c) conditioning seismic data for the nonlinear inversion.
Figure 4: Modeling error; a) examples of errors at specific noise levels; b) standard deviation of the homoscedastic error considered.
@@
$\sigma_{Fac.}$, $\sigma_{I_{P}}$ $(\frac{m}{s}\frac{g}{cm^{3}})$
@@
Data noise 1 0.05, 120 0.025, 0.5
Data noise 2 0.1, 230 0.05, 1
Data noise 3 0.2, 120 0.1, 2
@@ 3.2.1 Linear conditioning results
For the linear conditioning case, we considered 3 study cases contaminating $\bf{d_{obs}}$ (Figure 3b) with different uncorrelated Gaussian noise of absolute standard deviations ($\sigma_{Fac}$ and $\sigma_{I_{P}}$ in Table 2). Figure 5a shows the score magnitudes, expressed as L2-norms, for CDPS and DPS as a function of denoising/generative steps. For all settings, the scores of the data distribution have very similar values and decreasing trends; an average for the three noise levels is plotted as dashed lines. The likelihood scores show generally an inverse correlation between their magnitude and the data noise level, highlighting a stronger constraining effect of low-noise $\bf{d_{obs}}$. This correlation is negligible for the CDPS after 300 diffusion steps, with $\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{d}|\textbf{x}_{t})$ being approximately 30 times lower than $\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{x}_{t})$. A similar trend is observed in DPS only for the the largest data noise case (Data noise 3), while the likelihood score values generally increase with decreasing data error.
@@
For each scenario, we generated 100 conditional samples. Their convergence to $\bf{d_{obs}}$ is shown in Figure 5d as WRMSE between the predicted $\hat{\textbf{x}}_{0}^{(t)}$ ($\hat{\bf{x}}_{0}^{(t)}=\textbf{x}_{0}$ at $t=0$) and the conditional noisy $\bf{d_{obs}}$. Between 93 and 95% of the generated CDPS samples converged to a WRMSE $<1.1$. For DPS, none of the samples converged when considering the lowest noise levels, and 99% of the samples converged to a WRMSE <1.1 for the two larger-error cases. Overall, the WRMSE of the DPS samples approach lower values faster than CDPS during the denoising process; for Data noise 1, the DPS samples ($\hat{\textbf{x}}_{0}^{(t)}$) initially overfit the data noise (WRMSE < 1) and then divergence (WRMSE $\approx 1.5$). CDPS WRMSE values also show comparatively larger variance (i.e., DPS samples follow similar paths).
@@
Figure 5: Evolution of prior (dashed) and likelihood (continuous) score functions generative steps, expressed as L2-norm, for (a) linear conditioning and (b) nonlinear conditioning case studies; and WRMSE of 100 samples generated as solutions of the linear (c) and nonlinear (d) inverse problems. The values of magnitude of the three data noise assumed are indicated in 2.
@@
Figure 6a and b show the ensembles of realizations sampled by CDPS and DPS, respectively, for Data noise 2 case. The realizations’ generally converge correctly at well locations, and show increasingly larger variability with distance from the wells, towards the prior spatial uniformity. Considering, as example, the unconditioned leftmost area between x = 0 m and x = 20 m, CDPS better represents this uniformity (with p(sand) = 0.36 and $I_{P}$ = 7400 ±1080 m/s g/cm3), while DPS shows non-uniform channels distributions (p(sand) = 0.50±0.44 and $I_{P}$ = 7120 ±880  m/s g/cm3). We further attempted to sample the same posterior using pCN MCMC, considering $\beta$ = 0.03 and three independent chains, for 50000 iterations. After a burn-in time of 25000 iterations, only 60 parameters of the 120 conditioned parameters satisfied the $\hat{R}$ statistics threshold of 1.2. The sampled solutions’ ensembles are shown in Fig. 6c. Here, local conditioning is effective, but large variance (particularly evident from the facies distribution) highlight the points of non-convergence of the pCN. The unconditioned area between x = 0 m and x = 20 m, has p(sand) = 0.28 and $I_{P}$ = 7900 ±1030 m/s g/cm3 and shows facies and $I_{P}$ spatial patterns closer to those of the CDPS. However, the samples show poor mixing: on average, the converged parameters show autocorrelation values lower than 0.2 only after 400 samples (i.e., each chain has approximately 62 independent samples).
@@
Figure 6: Point-wise average and standard distribution of Facies and $I_{P}$ posterior samples from a) CDPS, b) DPS, c) pCN MCMC.
@@
Figure 7 shows the realizations of $I_{P}$ sampled by the three methods at the well locations, while their agreement with the target data ($\bf{d_{obs}}$ without noise) is summarized in Table 3. The DPS results show a comparatively poorer fit to the data than CDPS (Fig. 7b), showing larger WRMSE and logS values than for CDPS. While the pCN results fit the data within the noise assumptions (Fig. 7c), it shows larger logS and RMSE values than DPS and CDPS to Well 2 data. These discrepancies may be due to the lack of sufficient posterior samples as indicated by the limited convergence of the MCMC chains.
@@
WRMSE 0.70 0.74 0.69 0.79 0.78 0.86
LogS 6.3 7.1 6.3 6.5 6.7 6.4
LogS (Prior) 9.02 8.79
Figure 7: $I_{P}$ sampled distribution from (a) CDPS, (b) DPS, and (c) pCN MCMC, compared to the expected conditional data distribution $\bf{d_{obs}}$±$\pm\sigma_{I_{P}}$.
@@ 3.2.2 Nonlinear inversion results
We ran the nonlinear inversion example for three noise-levels in $\bf{d_{obs}}$ (Fig. 3c), considering absolute and relative components of the standard deviations ($\sigma_{c}$ and $r$, respectively, in Table 2). Similar to the linear case, the prior scores (Fig. 5b) display a decreasing trend that does not change significantly for different noise levels. The $\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{d}|\textbf{x}_{t})$ values of CDPS show a similar decreasing trend for the three noise levels, ranging from one-tenth to those of the $\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{x}_{t})$. Unlike the linear case, these scores show permanent differences between noise scales, with a maximum difference of one order of magnitude between the minimum and maximum data noise used. For DPS, the $\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{d}|\textbf{x}_{t})$ values are comparatively larger than CDPS and rapidly exceed the prior ones, becoming up to ten times larger (Data noise 1).
@@
The CDPS shows superior prediction accuracy. For the whole study area, the generated solutions’ ensemble capture the true $I_{P}$ much better than DPS, as shown by the lower values of logS in Table 4 and in the trace example in Figure 8h (x=25 m). Moreover, each realization reproduces the target $I_{P}$ values distribution with low $KL$ divergence (Table 4) and presents low-scale spatial variability patterns, which appear smoothed in DPS realizations (Fig. 8g).
In an attempt to compare the CDPS results with a high-quality posterior estimate, we ran the pCN method with $\beta$-values of 0.1, 0.05, 0.03, and 0.01, using three parallel chains. None of the runs converged towards a posterior distribution after 50000 iterations, indicating that the inversion problem is hard, and implicitly suggesting that the CDPS method performs well in practice.
@@
[…]
3.3 Computational performances
In this study, we did not consider any parallelization (see section 2.5). Moreover, we adopt a 2nd Heun solver when generating DM samples, and do not evaluate the costs/benefits of 1st solvers (e.g., Euler). The U-Net adopted in the DM ($1.31\times 10^{8}$ trainable parameters) required a total training time of 116 hours (600 epochs); the GAN ($1\times 10^{7}$ parameters) and VAE ($3.2\times 10^{7}$ parameters) used for comparison took 21 and 10 hours, respectively, (2000 epochs). The DM exceeded, however, the modeling accuracy of the fully trained GAN and VAE already at epoch 100, that is, after 19.5 hours of training (Appendix A).
With EDM, accurate sampling could be obtained using at least 18 diffusion steps, with no significant improvements when considering additional steps. The average generation time of the DM is 3 seconds multivariate realization (5 to 7 times longer than for GAN and VAE). For the linear conditioning case study (Section 3.2.1), the CDPS and DPS have similar computational time, of 0.6 seconds per time step in our tests. The relative time increase, compared to unconditional modeling, is mostly due to the backpropagation computation for the likelihood scores. For the non-linear case (Section 3.2.2), the CDPS had a larger computational cost, requiring 1.5 seconds per denoising step, mostly due to the additional Jacobian matrix computations. After testing multiple numbers of time-steps (from 18 to 2000 steps), we observed that both CDPS and DPS achieved their best modeling performances using 32 denoising steps. On the other hand we observed significant differences in inversion accuracy between CDPS and DPS, for nonlinear conditioning, as summarized in Table 5. Here, the CDPS had a near-optimal performance with 250 steps, while the DPS required a minimum of 1000 steps, which, in our tests, resulted in a significant reduction of the sampling time (approximately half), despite the increased computational costs of CDPS.
The time required by pCN MCMC for the linear case was substantially larger: the burn-in was reached after more than three days of computation, while the poor mixing of the sampled solutions show that much longer sampling would be necessary to recover the full posterior. We do not exclude that better hyperparameters would improve these performances. The application of pCN on the nonlinear case did not converge in any of the considered case.

### v2

@@ 2.5 Code implementation
The methods were implemented in Python (Python 3.12) using PyTorch (PyTorch 2.5) libraries with CUDA implementation. The training of all the generative models and the corresponding unconditional and conditional modeling were performed using a single GPU NVIDIA GeForce® GTX TITAN X (CUDA 12.2). The pCN sampling was performed on the same machine, generating DM samples using this GPU. No parallelization on different GPUs was considered. Training, modeling and sampling times mentioned in Section 3.3 are specific to this machine.
@@ 3 Results
We evaluate the generative performance of the trained DM by parameterizing a bivariate prior pdf, describing a sequence of lenticular sandy channels in a shale background (facies), and the associated acoustic impedances ($I_{P}$). We consider an area of 100 $\times$ 80 cells, assumed to have sides of 1 m (Fig. 1). The TI consists of 3000 samples (Fig. 1a) of MPS facies realizations, generated using Petrel software (SLB) based on a 3D conceptual geological model. Shales and sands are categorized as 0 and 1, respectively. The facies distribution ensemble is spatially uniform, with each cell having average p(sand) = 0.3 (Fig. 1c). The facies-dependent $I_{P}$ is represented using direct sequential co-simulations with multi-local distribution functions (Nunes et al., 2017; Soares, 2001). We account for facies-dependent $I_{P}$ marginal distributions (Fig. 1b) and spatial models (2D variograms; detailed in Appendix A). The average spatial distribution of $I_{P}$ (Fig. 1d) is primarily dependent on facies with an area of relatively lower values in the lower-right section of the considered area. The DM was trained for a total of 600 epochs (225000 steps). The GAN and VAE used for benchmarking were trained on the same TI, for a total of 2000 epochs, using the hyperparameters proposed by Miele et al. (2024).
Figure 1: Prior distribution as represented in the training dataset: (a) Facies and $I_{P}$ realizations; (b) $I_{P}$ distribution per facies; (c) probability of sand distribution; (d) point-wise average and standard deviation of $I_{P}$.
@@ 3.1 Unconditonal modeling
Prior modeling performances are assessed on 1000 unconditional samples generated from the trained DM, GAN and VAE each. The channel morphologies and $I_{P}$ spatial patterns reproduced by DM and GAN (Samples in Fig. 2a and c) are visually indistinguishable from those of the TI (Fig. 1a). For DM, the p(sand) and average $I_{P}$ match the prior accurately; Table 1 shows the average relative error for the two property types. Of the two benchmarking models, the GAN’s p(sand) and average $I_{P}$ distributions (Fig. 2c) show significant deviations from the prior. In fact, the overestimation of p(sand) in the central area may indicate mode collapse. Despite the smooth spatial features in the generated samples (Fig. 2e), the VAE performs better than GAN at capturing both properties average distributions, and displays even a larger accuracy than DM for facies (Table 1).
The joint distributions of facies and co-located $I_{P}$ obtained from the sampled realizations (Fig. 2b, d, and f) show significantly superior modeling ability of the DM (Fig. 2b). The facies classes are modeled with negligible errors (i.e., facies values different from 0 and 1); the corresponding $I_{P}$ marginal distributions (plotted along the y-axes) match accurately the prior means and standard deviations, fitting the data with a $KL$ divergence approximately five times lower than for GAN and VAE (Table 1). Further analysis of the networks unconditional modeling performances focusing on geostatistical metrics, are provided in Appendix A.
Table 1: Metrics for unconditional prior modeling using the trained DM and benchmarking GAN and VAE. The p(sand) and Avg. $I_{P}$ refer to the average models obtained from the realizations ensembles, shown in Figure 2a, c, e. KLKL: Kullback–Leibler divergence; Avg.: average; St. dev.: standard deviation.
@@
$7.0\%$
$1.2\%$
$6.87\pm 0.76$
$8.62\pm 0.69$
$1.3$
$1.5$
Figure 2: Summary of unconditional modeling performances of (a, b) DM, (c, d) GAN, and (e, f) VAE. Subplots (a), (c), (e): realizations and average distribution of facies and $I_{P}$ for (a) DM, (c) GAN and (e) VAE; subplots (b), (d), (f): joint and $I_{P}$ distributions for (b) DM, (d) GAN and (f) VAE.
@@ 3.2 Inverse modeling using CDPS
We evaluated the proposed CDPS by considering both linear and nonlinear inverse problems using a test realization from the prior distribution as our "True" facies and $I_{P}$ (Fig. 3a). The linear test case consists in generating samples that are locally conditioned on direct observations. The data, $\bf{d_{obs}}$, are obtained from two wells ("Well 1" and "Well 2" in Figure 3a). This problem is analogous to an image inpainting task in computer vision, where missing parts of images are generated conditioned on the known regions. The nonlinear example is a geophysical inversion problem with fullstack seismic data (Fig. 3c), assumed to be derived from the true $I_{P}$, and modeled as follows. First, vertical sequences of impedance contrasts ($R$) are computed following $R_{i}=(I_{P_{i+1}}-I_{P_{i}})/(I_{P_{i+1}}+I_{P_{i}})$, where the subscript $i$ indicates an $I_{P}$ sample. Then, each trace is convolved with a known seismic source wavelet; in our case, a Ricker wavelet with central frequency of 25Hz and resolution of 3 ms. For both case studies, we evaluated the CDPS under different Gaussian data noise conditions (Table 2).
@@
Using 300 validation examples from the TI we evaluated the modeling error $\sigma_{\hat{\textbf{x}}_{0}}^{(t)}$ for 1000 denoising steps. Examples of pixel-dependent uncertainties are shown in Fig. 4a, while the full homoscedastic noise sequence used is shown in Fig. 4b. Here, even if the facies classes are categorical, they are modeled as a continuous variable by DM and treated as such in the likelihood score computation. In Appendix B, we present applications of CDPS using well and seismic data simultaneously.
Figure 3: Test model used in inverse modeling: (a) real subsurface distribution ("True") and location of the two wells used for the linear inversion; (b) conditioning well log data for the linear inversion; (c) conditioning seismic data for the nonlinear inversion.
Figure 4: Modeling error: (a) examples of errors at specific noise levels; (b) standard deviation of the homoscedastic error considered. The values refer to the raw network output, of range $[0,1]$.
@@
$\sigma_{Fac.}$, $\sigma_{I_{P}}$ ($10^{6}\,\mathrm{Pa\cdot s/m}$)
@@
Data noise 1 $0.05,0.12$ $0.025,0.5$
Data noise 2 $0.1,0.23$ $0.05,1$
Data noise 3 $0.2,0.5$ $0.1,2$
@@ 3.2.1 Linear conditioning results
For the linear conditioning case, we considered three test cases involving the contamination of $\bf{d_{obs}}$ (Figure 3b) with different levels of uncorrelated Gaussian noise ($\sigma_{Fac}$ and $\sigma_{I_{P}}$, see Table 2). Figure 5a shows the score magnitudes, expressed as L2-norms, for CDPS and DPS as a function of denoising/generative steps. For all settings, the scores of the data distribution have very similar values and decreasing trends; an average for the three noise levels is plotted as dashed lines. The likelihood scores show generally an inverse correlation between their magnitude and the data noise level, highlighting a stronger constraining effect of low-noise $\bf{d_{obs}}$. This correlation is negligible for the CDPS after 300 diffusion steps, with $\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{d}|\textbf{x}_{t})$ being approximately 30 times lower than $\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{x}_{t})$. A similar trend is observed in DPS only for the the largest data noise case (Data noise 3), while the likelihood score values generally increase with decreasing data error.
@@
For each scenario, we generated 100 conditional samples. The convergence of the CDPS process is shown in Figure 5d as the WRMSE between the predicted $\hat{\textbf{x}}_{0}^{(t)}$ ($\hat{\bf{x}}_{0}^{(t)}=\textbf{x}_{0}$ at $t=0$) and $\bf{d_{obs}}$. Between 93 and 95% of the generated CDPS samples converged to a WRMSE <1.1<$1.$ For DPS, none of the samples converged when considering the lowest noise levels, and 99% of the samples converged to a WRMSE <1.1 for the two larger-error cases. Overall, the WRMSE of the DPS samples approach lower values faster than CDPS during the denoising process; for Data noise 1, the DPS samples ($\hat{\textbf{x}}_{0}^{(t)}$) initially overfit the data noise (WRMSE < 1) before diverging (WRMSE $\approx 1.5$). The WRMSE values obtained by CDPS also show comparatively larger variance (i.e., DPS samples follow similar paths).
@@
Figure 5: Evolution of prior (dashed) and likelihood (continuous) score functions generative steps, expressed as L2-norm, for (a) linear conditioning and (b) nonlinear conditioning case studies; and WRMSE of 100 samples generated as solutions of the (c) linear and (d) nonlinear inverse problems. The values of magnitude of the three data noise assumed are indicated in Table 2.
@@
Figures 6a and b show the ensembles of realizations sampled by CDPS and DPS, respectively, for the Data noise 2 case. The realizations generally converge correctly at well locations, and show increasingly larger variability with an increasing distance from the wells, approaching the prior model. Considering, as example, the unconditioned leftmost area between x = 0 m and x = 20 m, CDPS better represents this uniformity, with p(sand) = $0.36$ and $I_{P}=(7.40\pm 1.08)\times 10^{6}\,\mathrm{Pa\cdot s/m}$, while DPS shows non-uniform channel distributions, with p(sand) = $0.50$ and $I_{P}=(7.12\pm 0.88)\times 10^{6}\,\mathrm{Pa\cdot s/m}$. We further attempted to sample the posterior using pCN MCMC, considering $\beta$ = 0.03 and three independent chains, for 50000 iterations. After a burn-in time of 25000 iterations, only 60 parameters of the 120 conditioned parameters satisfied the $\hat{R}$ statistics threshold of 1.2. The ensemble of the sampled solutions are shown in Fig. 6c. Here, local conditioning is effective, but large variance (particularly evident in the facies distribution) highlight the points of non-convergence of the pCN. The unconditioned area between x = 0 m and x = 20 m has p(sand) = $0.28$ and $I_{P}=(7.90\pm 1.03)\times 10^{6}\,\mathrm{Pa\cdot s/m}$ and shows facies and $I_{P}$ spatial patterns closer to those of the CDPS. However, the samples show poor mixing: on average, the converged parameters show autocorrelation values lower than 0.2 only after 400 samples (i.e., each chain has approximately 62 independent samples).
@@
Figure 6: Point-wise average and standard distribution of Facies and $I_{P}$ posterior samples from (a) CDPS, (b) DPS, (c) pCN MCMC.
@@
Figure 7 shows the realizations of $I_{P}$ sampled by the three methods at the well locations, while their agreement with the target data ($\bf{d_{obs}}$ without noise) is summarized in Table 3. The DPS results present a comparatively poorer fit to the data than CDPS (Fig. 7b), showing larger WRMSE and logS values than for CDPS. While the pCN results fit the data within the noise assumptions (Fig. 7c), it shows relatively larger logS and RMSE values than CDPS to Well 2 data. These discrepancies may be due to the lack of sufficient posterior samples as indicated by the limited convergence of the MCMC chains.
@@
WRMSE $0.70$ $0.74$ $0.72$ $0.79$ $0.78$ $0.86$
LogS $13.2$ $14.0$ $13.2$ $13.3$ $13.6$ $13.7$
LogS (Prior) 15.9 15.7
Figure 7: $I_{P}$ sampled distribution from (a) CDPS, (b) DPS, and (c) pCN MCMC, compared to the expected conditional data distribution $\bf{d_{obs}}$$\pm\sigma_{I_{P}}$.
@@ 3.2.2 Nonlinear inversion results
We ran the nonlinear inversion example for three noise-levels in $\bf{d_{obs}}$ (Fig. 3c), considering absolute and relative components of the standard deviations ($\sigma_{c}$ and $r$, respectively, in Table 2). Similar to the linear case, the prior scores (Fig. 5b) display a decreasing trend that does not change significantly for different noise levels. The $\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{d}|\textbf{x}_{t})$ values of CDPS show a similar decreasing trend for the three noise levels, ranging from one-tenth to similar values as $\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{x}_{t})$. Unlike the linear case, these scores show persistent differences between noise scales, with a maximum difference of one order of magnitude between the minimum and maximum data noise used. For DPS, the $\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{d}|\textbf{x}_{t})$ values are comparatively larger than CDPS and rapidly exceed the prior scores, becoming up to ten times larger (Data noise 1).
@@
The CDPS has superior prediction accuracy. For the whole study area, the generated ensemble captures the true $I_{P}$ much better than DPS, as shown by the lower values of logS in Table 4 and in the trace example in Figure 8h (x = 25 m). Moreover, each realization reproduces the target $I_{P}$-distribution with a low $KL$ divergence (Table 4) and presents fine-scale spatial variability patterns, which appear smoothed in DPS realizations (Fig. 8g).
In an attempt to compare the CDPS results with a high-quality posterior estimate, we ran the pCN method with $\beta$-values of 0.1, 0.05, 0.03, and 0.01, using three parallel chains. None of the runs converged to the posterior distribution after 50000 iterations, indicating that the inversion problem is hard, and implicitly suggesting that the CDPS method performs well in practice.
@@
$KL_{I_{P}}$ $(10^{-3})$ $0.5\pm 0.2$ $1.3\pm 0.2$
3.3 Computational performance
In this study, we did not consider any parallelization (see section 2.5). Moreover, we adopt a 2nd order Heun solver when generating DM samples, and do not evaluate the costs/benefits of using a 1st order solvers (e.g., Euler). The U-Net adopted in the DM ($1.31\times 10^{8}$ trainable parameters) required a total training time of 116 hours (600 epochs); the GAN ($1\times 10^{7}$ parameters) and VAE ($3.2\times 10^{7}$ parameters) took 21 and 10 hours, respectively, to train (2000 epochs). The DM exceeded the modeling accuracy of the fully trained GAN and VAE already at epoch 100, that is, after 19.5 hours of training (see Appendix A).
With EDM, accurate sampling could be obtained when using at least 18 diffusion steps, with no significant improvements observed when considering additional steps. The average generation time of the DM is 3 seconds per multivariate realization (5 to 7 times longer than for GAN and VAE). For the linear conditioning case study (Section 3.2.1), the CDPS and DPS have similar computational time, 0.6 seconds per time step in our tests. The relative time increase, compared to unconditional modeling, is mostly due to the backpropagation computation for the likelihood scores. For the non-linear case (Section 3.2.2), the CDPS had a larger computational cost, requiring 1.5 seconds per denoising step, mostly due to the additional Jacobian matrix computations. After testing different numbers of time-steps (from 18 to 2000 steps), we observed that both CDPS and DPS achieved their best modeling performances using 32 denoising steps. On the other hand, we observed significant differences when considering inversion accuracy for the non-linear case, as summarized in Table 5. Here, the CDPS had a near-optimal performance using 250 steps, while the DPS required a minimum of 1000 steps, which, in our tests, resulted in a significant reduction of the sampling time (approximately half) when considering CDPS, despite its increased computational costs.
The time required by pCN MCMC for the linear case was substantially larger: the burn-in was reached after more than three days of computation, while the poor mixing of the sampled solutions show that much longer sampling would be necessary to recover the full posterior. We do not exclude that other hyperparameter values would improve the performance. The application of pCN on the nonlinear case did not converge in any of the considered cases.